OpenStudy (anonymous):

"Solve the system for X and Y" Y=-4x+21 2x+Y=9

OpenStudy (paxpolaris):

Using substitution: from the 1st equation you know \(y=-4x+21\) so where you see \(y\) in the 2nd equation you can replace it with \(-4x+21)\) you get:\[\large{2x+\left( -4x+21 \right)=9}\]\[\implies{2x-4x+21=9}\]